# MSCPT: Few-shot Whole Slide Image Classification with Multi-scale and Context-focused Prompt Tuning

### v2

@@ 3.3 Multi-scale Visual Descriptions Construction
In this part, we aim to generate pathological visual descriptions as pathological language prior knowledge to guide the hierarchical prompt tuning and instances aggregation. To reduce manual workload, large language models (LLMs) are employed to generate descriptions related to different diseases. That is, we enter the question “We are studying Cancer Category. Please list $\textbf{C}^{l}$ visual descriptions at 5×5$\times$5 × magnification and $\textbf{C}^{h}$ visual descriptions at 20×20$\times$20 × magnification observed in H&E-stained histological images of Cancer Sub-category.” into the LLM. And then we can get the multi-scale visual description sets $T^{low}=\{T^{low}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{l}\}$ and $T^{high}=\{T^{high}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{h}\}$. $K$ represents the number of WSI categories, and $C^{l}$ and $C^{h}$ denote the counts of low-level and high-level descriptions, respectively.
@@ 3.4.2 Hierarchical Prompt Tuning
[…]
@@ 3.6 Non-Parametric Cross-Guided Pooling
$Logits^{high}&=h_{\text{topK}}\left(\tilde{\mathbf{P}}^{high}\cdot{\mathbf{Z}^{high}}^{\mathrm{T}}\right)\\ &\quad+h_{\text{topK}}\left(\tilde{\mathbf{P}}^{high}\cdot{\mathbf{Z}^{low}}^{\mathrm{T}}\right),\end{aligned}$ (11)
$Logits^{low}&=h_{\text{topK}}\left(\tilde{\mathbf{P}}^{low}\cdot{\mathbf{Z}^{low}}^{\mathrm{T}}\right)\\ &\quad+h_{\text{topK}}\left(\tilde{\mathbf{P}}^{low}\cdot{\mathbf{Z}^{high}}^{\mathrm{T}}\right),\end{aligned}$ (12)
[…]
@@
{tblr}
colsep=1.5pt, rowsep = 0.05cm, column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, column11 = c, column12 = c, column13 = c, column14 = c, column15 = c, column16 = c, column17 = c, column18 = c, column19 = c, column20 = c, column21 = c, column22 = c, column23 = c, row16 = bg=lightgray, row30 = bg=lightgray, row44 = bg=lightgray, cell11 = c=2,r=2c, cell13 = r=2, cell14 = c=3, cell18 = c=3, cell112 = c=3, cell116 = c=3, cell120 = c=3, cell31 = r=14c, bg=lightgray, cell171 = r=14c, bg=lightgray, cell311 = r=14c, bg=lightgray, hline1,17,31,45 = -0.08em, hline2 = 4-6,8-10,12-14,16-18,20-22, hline3 = -, Methods & Train. Param TCGA-NSCLC
@@
CLIP Pretrained on 400M Natural Image Pairs Max-pooling 197K 63.80±6.84 60.40±4.76 60.70±4.75 84.51±3.21 65.83±2.72 69.26±2.33 87.76±1.45 62.36±5.63 63.58±4.05 59.74±1.88 57.03±1.71 63.85±2.06 72.12±1.71 51.68±3.55 56.48±1.50
@@
PLIP Pretrained on 208K Pathology Image Pairs Max-pooling 197K 71.78±4.13 66.40±3.51 66.66±3.42 95.18±0.63 81.63±0.92 84.30±1.30 91.57±1.35 71.06±3.30 71.09±3.39 65.83±2.66 59.92±1.62 66.12±3.57 78.72±3.05 59.24±3.22 62.86±3.13
@@
CONCH Pretrained on 1.17M Pathology Image Pairs Max-pooling 131K 94.68±1.91 87.97±1.85 87.98±1.85 97.42±0.25 90.20±0.91 91.35±0.65 95.18±0.56 80.78±1.77 80.99±1.87 76.00±2.11 67.18±2.93 71.38±3.17 84.42±2.57 65.90±3.02 69.85±3.62
@@ 4.1.1 Datasets
TCGA-NSCLS, TCGA-RCC, and UBC-OCEAN datasets are used to examine the cancer subtyping ability of MSCPT. TCGA-NSCLC contains 1041 non-small cell lung cancer WSIs, comprising 530 lung adenocarcinoma (LUAD) and 511 lung squamous cell carcinoma (LUSC) WSIs. TCGA-RCC includes 873 renal cell carcinoma WSIs: 121 chromophobe renal cell carcinoma (CHRCC), 455 clear-cell renal cell carcinoma (CCRCC), and 297 papillary renal cell carcinoma (PRCC). UBC-OCEAN comprises 538 ovarian WSIs, with 99 clear cell carcinoma (CC), 124 endometrioid carcinoma (EC), 221 high-grade serous carcinoma (HGSC), 47 low-grade serous carcinoma (LGSC), and 46 mucinous carcinoma (MC) WSIs.
@@ 4.1.4 Implementation Details
Following CLAM [28], the original WSIs were initially processed using the Otsu thresholding algorithm to remove the background parts. Subsequently, the WSIs were segmented into multiple non-overlapping patches of $256\times 256$ pixels at 5×5$\times$5 × and 20×20$\times$20 × magnification levels. We evaluate MSCPT on CLIP [12], PLIP [19], and CONCH [20], with the first two using ViT-B/32 [38] and CONCH using ViT-B/16 [38] as their visual towers. Apart from MSCPT, Metaprompt, and DSMIL, which used both 5×5$\times$5 × and 20×20$\times$20 × magnification patches, the remaining methods solely relied on 20×20$\times$20 × magnification patches as inputs. To ensure fair comparisons across different shot settings, 80% of each dataset was used for testing, while $K$ samples per class ($K=16,8,4,2$) were drawn from the remaining 20% for training.
All methods used the Adam optimizer (learning rate: $41e-$ 4, weight decay: $51e-$ 5) with a batch size of 1. Training was fixed at 100 epochs for CLIP and 50 for PLIP and CONCH, with early stopping applied. We chose GPT-4 [43] to generate pathological visual descriptions, providing 10 low-level and 30 high-level visual descriptions for each category of WSIs (i.e., $C^{l}=10$ and $C^{h}=30$). For MSCPT and Metaprompt, we selected 30 patches for each category at 5×5$\times$5 × magnification for visual prompt tuning. The lengths of the global prompts $p_{glob}$ in both image and text encoders were uniformly set to 2. All experiments were conducted on a workstation with eight NVIDIA A800 GPUs.
@@
{tblr}
colsep=1.5pt, rowsep = 0.05cm, column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, column11 = c, column12 = c, column13 = c, column14 = c, column15 = c, column16 = c, column17 = c, column18 = c, column19 = c, column20 = c, column21 = c, column22 = c, column23 = c, row16 = bg=lightgray, row30 = bg=lightgray, row44 = bg=lightgray, cell11 = c=2,r=2c, cell13 = c=3, cell17 = c=3, cell111 = c=3, cell115 = c=3, cell119 = c=3, cell31 = r=14c, bg=lightgray, cell171 = r=14c, bg=lightgray, cell311 = r=14c, bg=lightgray, hline1,17,31,45 = -0.08em, hline2 = 3-5,7-9,11-13,15-17,19-21, hline3 = -, Methods & TCGA-NSCLC
(Subtyping, 2 classes) ~ TCGA-RCC
(Subtyping, 3 classes) ~ UBC-OCEAN
(Subtyping, 5 classes) ~ TCGA-BRCA
(Recurrence Prediction, 2 classes) ~ PANDA
@@
2-shot 2 Training Samples per Class Max-pooling 80.85±5.19 64.68±6.49 67.31±4.22 86.62±5.10 69.08±8.05 73.21±7.63 84.82±2.76 58.70±5.87 62.65±5.75 67.92±6.34 61.66±7.31 67.60±5.07 70.31±3.67 51.20±4.20 56.44±3.67
@@
4-shot 4 Training Samples per Class Max-pooling 84.75±4.38 76.52±5.40 76.66±5.38 92.80±2.37 77.72±5.20 80.89±4.58 90.18±1.52 68.34±6.33 69.85±4.20 72.55±6.54 64.69±4.66 68.95±2.82 75.09±2.06 56.71±2.99 61.33±3.18
@@
8-shot 8 Training Samples per Class Max-pooling 91.54±1.81 83.41±3.23 83.46±3.16 97.39±0.54 88.47±1.53 90.17±1.10 91.72±1.15 71.43±1.55 73.84±2.08 74.62±1.91 66.56±1.69 69.70±2.10 77.49±4.34 59.38±4.25 64.19±4.15
@@ 4.2 Comparisons with State-of-the-Art
The experimental results under the 16-shot setting are displayed in Table 1. We observed some intriguing insights, such as complex and parameter-heavy methods like TransMIL and RRT-MIL underperformed despite their strong performance with full data training. Conversely, less parameterized methods such as ABMIL and CLAM exhibited slightly better performance. This is because traditional MIL-based methods require a large number of WSIs for training, and the more parameters they have, the more training data is needed. Furthermore, after adapting the prompt tuning methods designed for natural images (i.e., CoOp, CoCoOp, and Metaprompt) to tasks at the WSI-level, these methods outperform traditional MIL-based methods when based on CLIP and achieve comparable performance when using PLIP and CONCH. Relatively few parameters contribute to this result. Additionally, we found that Metaprompt outperforms CoOp across most metrics, thanks to its integration of visual prompt tuning and multi-scale information. This result motivates us to pursue visual prompt tuning and develop more effective multi-scale information integration modules. Despite prompt tuning methods designed for Few-shot Weakly Supervised WSI Classification (FSWC) tasks having a relatively higher number of parameters, they exhibit good performance. This leverages VLMs’ prior knowledge through visual and text prompt tuning, reducing the need for extensive training samples.
@@
Prior Knowl. Multi Scale MHPT ISGPT NPCGP TCGA-NSCLC (PLIP-based, 16-shot) PANDA (CONCH-based, 4-shot)
@@
Global High Low UBC-OCEAN (CONCH-based, 4-shot)
@@ 4.3 Results with Fewer Training Samples
To further validate MSCPT’s performance, we conducted experiments under 8-, 4-, and 2-shot settings. Notably, with limited training samples, sample selection critically influences model performance [6]. However, this influence diminishes as the performance of the VLM improves. Accordingly, this experiment was conducted based on CONCH, with the results presented in Table 2.
@@
Methods Train. Param. TCGA-RCC (CLIP-based, 16-shot)
@@ 4.4.4 Effects of Image-text Similarity-based Graph Prompt Tuning
Methods TCGA-NSCLC (PLIP-based, 16-shot)
@@ 4.4.5 Effects of Instance Aggregation
Methods TCGA-RCC (CLIP-based, 16-shot)

### v3

@@ 3.3 Multi-scale Visual Descriptions Construction
In this part, we aim to generate pathological visual descriptions as pathological language prior knowledge to guide the hierarchical prompt tuning and instances aggregation. To reduce manual workload, large language models (LLMs) are employed to generate descriptions related to different diseases. That is, we enter the question “We are studying Cancer Category. Please list $\textbf{C}^{l}$ visual descriptions at $5\times$ magnification and $\textbf{C}^{h}$ visual descriptions at $20\times$ magnification observed in H&E-stained histological images of Cancer Sub-category.” into the LLM. And then we can get the multi-scale visual description sets $T^{low}=\{T^{low}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{l}\}$ and $T^{high}=\{T^{high}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{h}\}$. $K$ represents the number of WSI categories, and $C^{l}$ and $C^{h}$ denote the counts of low-level and high-level descriptions, respectively.
@@ 3.4.2 Hierarchical Prompt Tuning
$\begin{aligned} \small\left[C^{i},\_,\_,p^{i}_{high},E^{i}\right]&=T_{i}\left[C^{i-1},p_{glob}^{i-1},p_{low}^{i-1},p_{high}^{i-1},E^{i-1}\right],\\ i&=1,2,3,...,L,\end{aligned}$ (4)
@@ 3.6 Non-Parametric Cross-Guided Pooling
$\small\begin{aligned} Logits^{high}&=h_{\text{topK}}\left(\tilde{\mathbf{P}}^{high}\cdot{\mathbf{Z}^{high}}^{\mathrm{T}}\right)\\ &\quad+h_{\text{topK}}\left(\tilde{\mathbf{P}}^{high}\cdot{\mathbf{Z}^{low}}^{\mathrm{T}}\right),\end{aligned}$ (11)
$\begin{aligned} Logits^{low}&=h_{\text{topK}}\left(\tilde{\mathbf{P}}^{low}\cdot{\mathbf{Z}^{low}}^{\mathrm{T}}\right)\\ &\quad+h_{\text{topK}}\left(\tilde{\mathbf{P}}^{low}\cdot{\mathbf{Z}^{high}}^{\mathrm{T}}\right),\end{aligned}$ (12)
$\begin{aligned} Logits^{overall}&=\frac{1}{2}\left(Logits^{high}+Logits^{low}\right).\end{aligned}$ (13)
@@
Methods Train. Param
@@
CLIP Pretrained on 400M Natural Image Pairs
Max-pooling 197K 63.80±6.84 60.40±4.76 60.70±4.75 84.51±3.21 65.83±2.72 69.26±2.33 87.76±1.45 62.36±5.63 63.58±4.05 59.74±1.88 57.03±1.71 63.85±2.06 72.12±1.71 51.68±3.55 56.48±1.50
@@
PLIP Pretrained on 208K Pathology Image Pairs
Max-pooling 197K 71.78±4.13 66.40±3.51 66.66±3.42 95.18±0.63 81.63±0.92 84.30±1.30 91.57±1.35 71.06±3.30 71.09±3.39 65.83±2.66 59.92±1.62 66.12±3.57 78.72±3.05 59.24±3.22 62.86±3.13
@@
CONCH Pretrained on 1.17M Pathology Image Pairs
Max-pooling 131K 94.68±1.91 87.97±1.85 87.98±1.85 97.42±0.25 90.20±0.91 91.35±0.65 95.18±0.56 80.78±1.77 80.99±1.87 76.00±2.11 67.18±2.93 71.38±3.17 84.42±2.57 65.90±3.02 69.85±3.62
@@ 4.1.1 Datasets
TCGA-NSCLS, TCGA-RCC, and UBC-OCEAN datasets are used to examine the cancer subtyping ability of MSCPT. TCGA-NSCLC contains 1041 non-small cell lung cancer WSIs, comprising 530 lung adenocarcinoma (LUAD) and 511 lung squamous cell carcinoma (LUSC) WSIs. TCGA-RCC includes 873 renal cell carcinoma WSIs: 121 chromophobe renal cell carcinoma (CHRCC), 455 clear-cell renal cell carcinoma (CCRCC), and 297 papillary renal cell carcinoma (PRCC). UBC-OCEAN [41] comprises 538 ovarian WSIs, with 99 clear cell carcinoma (CC), 124 endometrioid carcinoma (EC), 221 high-grade serous carcinoma (HGSC), 47 low-grade serous carcinoma (LGSC), and 46 mucinous carcinoma (MC) WSIs.
@@ 4.1.4 Implementation Details
Following CLAM [28], the original WSIs were initially processed using the Otsu thresholding algorithm to remove the background parts. Subsequently, the WSIs were segmented into multiple non-overlapping patches of $256\times 256$ pixels at $5\times$ and $20\times$ magnification levels. We evaluate MSCPT on CLIP [12], PLIP [19], and CONCH [20], with the first two using ViT-B/32 [38] and CONCH using ViT-B/16 [38] as their visual towers. Apart from MSCPT, Metaprompt, and DSMIL, which used both $5\times$ and $20\times$ magnification patches, the remaining methods solely relied on $20\times$ magnification patches as inputs. To ensure fair comparisons across different shot settings, 80% of each dataset was used for testing, while $K$ samples per class ($K=16,8,4,2$) were drawn from the remaining 20% for training.
All methods used the Adam optimizer (learning rate: 1e41e-4, weight decay: 1e51e-5) with a batch size of 1. Training was fixed at 100 epochs for CLIP and 50 for PLIP and CONCH, with early stopping applied. We chose GPT-4 [44] to generate pathological visual descriptions, providing 10 low-level and 30 high-level visual descriptions for each category of WSIs (i.e., $C^{l}=10$ and $C^{h}=30$). For MSCPT and Metaprompt, we selected 30 patches for each category at $5\times$ magnification for visual prompt tuning. The lengths of the global prompts $p_{glob}$ in both image and text encoders were uniformly set to 2. All experiments were conducted on a workstation with eight NVIDIA A800 GPUs.
@@
Methods
(Subtyping, 2 classes) TCGA-RCC
(Subtyping, 3 classes) UBC-OCEAN
(Subtyping, 5 classes) TCGA-BRCA
(Recurrence Prediction, 2 classes) PANDA
@@ Methods
2-shot 2 Training Samples per Class
Max-pooling 80.85±5.19 64.68±6.49 67.31±4.22 86.62±5.10 69.08±8.05 73.21±7.63 84.82±2.76 58.70±5.87 62.65±5.75 67.92±6.34 61.66±7.31 67.60±5.07 70.31±3.67 51.20±4.20 56.44±3.67
@@
4-shot 4 Training Samples per Class
Max-pooling 84.75±4.38 76.52±5.40 76.66±5.38 92.80±2.37 77.72±5.20 80.89±4.58 90.18±1.52 68.34±6.33 69.85±4.20 72.55±6.54 64.69±4.66 68.95±2.82 75.09±2.06 56.71±2.99 61.33±3.18
@@
8-shot 8 Training Samples per Class
Max-pooling 91.54±1.81 83.41±3.23 83.46±3.16 97.39±0.54 88.47±1.53 90.17±1.10 91.72±1.15 71.43±1.55 73.84±2.08 74.62±1.91 66.56±1.69 69.70±2.10 77.49±4.34 59.38±4.25 64.19±4.15
@@ 4.2 Comparisons with State-of-the-Art
The experimental results under the 16-shot setting are displayed in Table 3.6. We observed some intriguing insights, such as complex and parameter-heavy methods like TransMIL and RRT-MIL underperformed despite their strong performance with full data training. Conversely, less parameterized methods such as ABMIL and CLAM exhibited slightly better performance. This is because traditional MIL-based methods require a large number of WSIs for training, and the more parameters they have, the more training data is needed. Furthermore, after adapting the prompt tuning methods designed for natural images (i.e., CoOp, CoCoOp, and Metaprompt) to tasks at the WSI-level, these methods outperform traditional MIL-based methods when based on CLIP and achieve comparable performance when using PLIP and CONCH. Relatively few parameters contribute to this result. Additionally, we found that Metaprompt outperforms CoOp across most metrics, thanks to its integration of visual prompt tuning and multi-scale information. This result motivates us to pursue visual prompt tuning and develop more effective multi-scale information integration modules. Despite prompt tuning methods designed for Few-shot Weakly Supervised WSI Classification (FSWC) tasks having a relatively higher number of parameters, they exhibit good performance. This leverages VLMs’ prior knowledge through visual and text prompt tuning, reducing the need for extensive training samples.
@@
Prior Knowl. Multi Scale MHPT ISGPT NPCGP
TCGA-NSCLC
(PLIP-based, 16-shot)
PANDA
(CONCH-based, 4-shot)
@@
Global High Low
UBC-OCEAN
(CONCH-based, 4-shot)
@@ 4.3 Results with Fewer Training Samples
To further validate MSCPT’s performance, we conducted experiments under 8-, 4-, and 2-shot settings. Notably, with limited training samples, sample selection critically influences model performance [6]. However, this influence diminishes as the performance of the VLM improves. Accordingly, this experiment was conducted based on CONCH, with the results presented in Table 4.1.4.
@@
Methods Train. Param.
TCGA-RCC
(CLIP-based, 16-shot)
@@ 4.4.4 Effects of Image-text Similarity-based Graph Prompt Tuning
Methods
TCGA-NSCLC
(PLIP-based, 16-shot)
@@ 4.4.5 Effects of Instance Aggregation
Methods
TCGA-RCC
(CLIP-based, 16-shot)